OpenStudy (anonymous):

the length of the rectangle is 3m less than 5 times the width. the area of the rectangle is 36m^2. find the dimensions of the rectangle.

OpenStudy (anonymous):

Okay let's come up with some variables. \(w\) is width, \(l\) is length. "the length of the rectangle is 3m less than 5 times the width" Create an equation from this!!! "the area of the rectangle is 36m^2." Ooooh, it's another equation. Make it!

OpenStudy (anonymous):

5w-3+w=36m^2

OpenStudy (anonymous):

Well, that's a good start... but do you add length and width or do you multiply them?

OpenStudy (anonymous):

to get the area.

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

Okay, then your equation should have multiplication.

OpenStudy (anonymous):

were?

OpenStudy (anonymous):

Hold on though... I think we're getting ahead of ourselves. First of all, what is the equation for this statement: ""the length of the rectangle is 3m less than 5 times the width"

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

??? Using \(l\) and \(w\) what is the equation?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

"the length of the rectangle is 3m less than 5 times the width" write this in algebra

OpenStudy (anonymous):

yes i know that was in my equation

OpenStudy (anonymous):

I know, but I just want to be on the same page here.

OpenStudy (anonymous):

When you do so many steps at once you make mistakes.

OpenStudy (anonymous):

There are two equations, but you only had one equation and it had a mistake.

OpenStudy (anonymous):

Which is why I just want to do thinks with simple steps first.

OpenStudy (anonymous):

`the length of the rectangle is 3m less than 5 times the width`\[ l = 5w-3 \] `the area of the rectangle is 36m^2`\[ lw = 36 \]

OpenStudy (anonymous):

Now when you do the substitution you get \[ (5w-3)(w)=36 \]

OpenStudy (anonymous):

ya i know

OpenStudy (anonymous):

If you know, then show me =)

OpenStudy (anonymous):

I cannot read minds unfortunately.

OpenStudy (anonymous):

oomg.............

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

this is taking way to long! all i needd is a equation then i can solve it but you just brought me where i started and needed help!!!!!!!

OpenStudy (anonymous):

No you had \[ 5w-3+w=36m^2 \]and it was wrong.

OpenStudy (anonymous):

You're not asking any questions either.

OpenStudy (anonymous):

I'm not going to do it for you. Sometimes I do most of it for a person, sometimes not. You didn't win that lottery.

OpenStudy (anonymous):

i really dont care anymore its eleven a clock and now i am pissy

OpenStudy (anonymous):

\[ (5w-3)(w)=5w^2-3w=36 \]You know how to solve quadratic equation?

OpenStudy (anonymous):

ya but this one doesnt had up

OpenStudy (anonymous):

Doesn't had up? \[ 5w^2-3w-36 = 0 \]Who needs up?\[ a=5,b=-3,c=-36 \]

OpenStudy (anonymous):

its 12 and 3

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=5w%5E2%E2%88%923w%E2%88%9236%3D0 I'm getting \[ w=3, w=-\frac{12}{5} \]Do you think width is negative?

OpenStudy (anonymous):

flutter me i hate life i am tired i need to eat

OpenStudy (anonymous):

i really dont know anymore

OpenStudy (anonymous):

Sorry, width is only positive. That means \(w=3\).

OpenStudy (anonymous):

\[ l = 5w-3= 5(3)-3 \]

OpenStudy (anonymous):

i reealyy hate life so bye :......(

OpenStudy (anonymous):

http://www.youtube.com/watch?v=ZQ7oqmikZDQ